Gary has a box of coins that he uses when playing Poker. The box contains 53 coins consisting of nickels, dimes and quarters. The number of quarters is equal to the number of dimes. The total value is $6.80. How many if each type of coin does Gary have in the box?
Answer by TimothyLamb(4379)

x + y + z = 53
5x + 10y + 25z = 680
y = z
---
put the system of linear equations into standard form
---
x + y + z = 53
5x + 10y + 25z = 680
y - z = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
the solver gives the solution to the linear system as:
x= 19.8
y= 16.6
z= 16.6
---
answer:
this problem statement contains an error.
the solution to a system of linear equations that represents numbers of coins cannot yield fractional numbers of coins.
in other words, there cannot be 0.8 nickels or 0.6 dimes.
this means that the problem statement is flawed.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php